OpenStudy (dreamsawakes):

If the sin 60° = square root of three over two, then which statement is true?

jhonyy9 (jhonyy9):

so i think that cos 60° = 1/2 will be true sure

OpenStudy (dreamsawakes):

cos 30° = square root of three over two, because the cosine and sine are complements cos 120° = square root of three over two, because the cosine and sine are supplements cos 30° = 0, because the cosine and sine are complements cos 120° = 0, because the cosine and sine are supplements

OpenStudy (dreamsawakes):

those are the choices

jhonyy9 (jhonyy9):

so than the first one is right bc. than sin60° = sqrt3 /2 so than cos 30° will be = sqrt3 /2 sure looke please this is the table with details sin 30 = 1/2 cos 30 = sqrt3/2 sin 45 = sqrt2 /2 cos 45 = sqrt2 /2 sin 60 = sqrt3 /2 cos 60 = 1/2 hope so much that thses will help you sure easy understanding it right good luck was my pleasure bye bye

OpenStudy (dreamsawakes):

Thank you

jhonyy9 (jhonyy9):

np was my pleasure bye bye and anytime than you wan and i m here online